Find equations in standard form of the lines through point P that are
A.) parallel to, and
B.) perpendicular to, line L.
P(0,3); L: x+y=5
Please help me. My teacher in school does not teach us, he tells us to read it and do it on our own and it seems simple but I'm not sure what to do.

Found 2 solutions by richwmiller, Alan3354:
Answer by richwmiller(17219)

Many of the tutors here are or were teachers.
Does it seem wise to complain about teachers?

Answer by Alan3354(69443)

I can work this problem. My being able to work it won't help you, tho.
You have to learn how to work these.
Do it like this one:
-----------------------
Write in standard form the eqation of a line that satisfies the given conditions. Perpendicular to 9x+3y=36, through (1,2)
-----------------
Find the slope of the line. Do that by putting the equation in slope-intercept form, y = mx + b. That means solve for y.
9x+3y = 36
3y= - 9x + 36
y = -3x + 13
The slope, m = -3
------------------
The slope of lines parallel have the same slope.
The slope of lines perpendicular is the negative inverse, m = +1/3
----------------
Use y = mx + b and the point (1,2) to find b.
2 = (1/3)*1 + b
b = 5/3
The equation is y = (1/3)x + 5/3 (slope-intercept form)
x - 3y = -5 (standard form)
------------------------
For further assistance, or to check your work, email me via the thank you note or at moralloophole@aol.com